OpenStudy (anonymous):

A customer has an offer for 5 portrait sheets for 9.99 with additional portrait sheets priced at 19.99 each. If the customer desires to purchase 10 portrait sheets, how much would she save if she instead purchases a 10 portrait sheet package for $99.99? $19.99 Not Here $19.95 $9.99 $9.95

OpenStudy (anonymous):

Offer for 5 plus five = 9.99 + 5(19.99) = 9.99 + 99.50 = 109.49 Compare ten at 99.99 So saving = 109.49 - 99.99 = 9.50 So 'not here'